# Block-Attention for Efficient RAG

### v2

@@ Abstract
We introduce Block-Attention, an attention mechanism designed to address the increased inference latency and cost in Retrieval-Augmented Generation (RAG) scenarios. Unlike existing works that encodes the whole context, its main idea lies in dividing the retrieved documents into blocks, where each block calculates key-value (KV) states independently except for the final block. In RAG scenarios, by defining each passage as a block, Block-Attention enables us to pre-compute the KV states for all passages and cache them in memory, significantly reducing the latency and the computation cost during inference. The implementation involves block segmentation, positional encoding calculation, and fine-tuning the LLM to adapt to the Block-Attention mechanism. Experiments on four RAG benchmarks demonstrate that after block fine-tuning, the Block Attention model can achieve performance comparable to (68.4% vs 67.9% on Llama3) or even better (62.8% vs 59.6% on Mistral) than self-attention models. Notably, Block-Attention reduces the TTFT (the time to first token) and FLOPs (floating point operations) to a very low level. It only takes 45 ms to output the first token for an input sequence with a total length of 32K. Compared with the self-attention model, the time consumption and corresponding FLOPs are reduced by 98.7% and 99.8%, respectively. [^1] Codes are available at: https://github.com/TemporaryLoRA/Block-Attention
@@ 1 Introduction
However, everything has two sides, and RAG is no exception. Generally, for each user query, in order to ensure that a passage with the “correct knowledge” is successfully recalled, we retrieve multiple passages (between 5 to 30 in most scenarios) Kwiatkowski et al. (2019); Joshi et al. (2017) and incorporate them into the input prompt of the LLM. Within this new and significantly longer input prompt, the inference latency, particularly the time to first token (TTFT), of a RAG-LLM will be much higher than that of a non-RAG-LLM Li et al. (2023); Zhu et al. (2024).
Given that the content of passages in the external databases is already determined, one might naturally consider the feasibility of caching them for fast inference. Nonetheless, for autoregressive LLMs, the KV states are inherently context-dependent, which means the KV states for the same passage will vary in different contexts. As a result, when encountering an unseen query, the model must undertake a complete re-encoding of the KV states to ensure the accuracy of its predictions.
In this paper, we propose the Block-Attention method, which reduces the TTFT of RAG-LLMs to a level nearly equivalent to that of non-RAG LLMs, while fully maintaining the same accuracy level. As shown in Figure 1, the main idea of Block-Attention is to divide the entire input sequence into several blocks. Each block independently calculates its KV states through self-attention, without any attention to other blocks. Only the final block is able to attend other blocks. When utilizing Block-Attention in RAG scenarios, we may achieve substantial benefits by defining each passage as a single block and caching its KV states in the memory. The Block-Attention method has effectively reduced the average TTFT in experiments. When the length of the input sequence is 32K and the length of the user input is 50, the Block-Attention method can achieve a 98.7% reduction in TTFT (from 3638 ms to 45 ms), along with a 99.8% reduction in FLOPs of first token.  [^2] Given that the KV cache is already a mature and low-cost technology (Qin et al., 2024), in this paper we do not take the cost of KV cache into account.
@@
The implementation of Block-Attention can be easily achieved through the following steps: 1) Encode all blocks except the last one separately; 2) calculating the positional encoding for each based on its position in the input prompt; 3) Integrating these blocks to compute the KV states for the final block. However, the primary challenge in utilizing Block-Attention is that the LLMs have not been exposed to such an attention mechanism during their training, leading to difficulties in accurately interpreting the input prompt. In our preliminary experiments, we attempted to directly implement Block-Attention in the LLMs without updating any parameters. Unfortunately, this approach resulted in a substantial decrease in performance, with the average accuracy of Llama3 on four RAG benchmarks falling from 67.9% to 48.0%.
@@
We evaluate Block-Attention on four RAG benchmarks. Experimental results demonstrate that after fine-tuning, the average accuracy of the Block-Attention model on the benchmarks can remain comparable to (68.4% vs 67.9% on Llama3) or even slightly exceed (62.8% vs 59.6% on Mistral) the performance of self-attention models. In terms of efficiency, we counted the TTFT of the block-attention model when the length of user input is 50 and the total length of the input sequence gradually increases. We found that the longer the total length, the more obvious the improvement of block-attention on efficiency. When the length of the input sequence reaches 32K, the TTFT of the block-attention model is only 45 ms, which is 1.3% of that of the self-attention model.
@@ 2.1 Main Idea
In the scenario of RAG, a given passage frequently recurs in the retrieval results for numerous queries, making it a natural consideration to explore the possibility of reusing its KV states. Unfortunately, the auto-regressive nature of the encoding approach, as previously outlined, imposes constraints that render the KV states nearly non-reusable across distinct queries. Our research is dedicated to examining a novel attention mechanism, Block-Attention, that designed to necessitate the re-computation of only the altered text blocks between two input sequences, thereby achieving an outcome equivalent to that of full sequence re-encoding.
@@ 2.3 Positional Encoding
2) we set $x_{i}$’s positional encoding to zero, which is equivalent to rotating it counterclockwise by $i\theta$ degrees
@@
3) Then, by performing a clockwise rotation of $(i_{\Delta})\theta$ degrees, we obtain the final positional encoding vector.
@@ 2.5 Inference
In inference, the Block-Attention model retrieves the KV states of blocks from the KV cache and concatenates them into the complete input KV states. The detailed process of the inference stage is depicted in Figure 2. Initially, we query and extract the KV states of the first $k-1$ blocks from the cache. Then, based on the position of each block within the input sequence, we calculate their positional encoding using Equation 3. Finally, using the KV states of the first $k-1$ blocks, we compute the KV states of the last block as well as the model output.
@@ 3 Experiments
After the above analysis, most readers may have the following two concerns about Block-Attention: 1) Can the Block-Attention model achieve the same accuracy as self-attention in RAG scenarios? 2) How much can the Block-Attention mechanism improve the RAG efficiency? The following experimental results will reveal the answers to these two questions for everyone. In Sections 3.1, we analyze the accuracy of Block-Attention models. Meanwhile, in Section 3.2, we demonstrate the efficiency of Block-Attention.
Datasets
We evaluate the performance of the Block-Attention method on four RAG benchmarks: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), HotpotQA (HQA) (Yang et al., 2018), and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020). Specifically, we employ the training datasets of TQA and 2Wiki for fine-tuning models, followed by evaluations on all four respective test sets.
In training, we have randomly sampled 20,000 instances from both the TQA and 2wiki datasets to form our training set. The model’s input comprises (1) the question and (2) 10 retrieved passages. To obtain these 10 passages, we employed the Contriver [^3] https://github.com/facebookresearch/contriever to identify the 10 most relevant passages from the provided dataset. Additionally, due to the fact that the original answers in the TQA and 2wiki datasets may not be included in the retrieved passages, this could lead the model to overlook the content of the retrieved passages and generate outputs directly. Instead, we utilize the outputs from GPT-4 as the model’s output to ensure consistency between the retrieved passages and the final output.
@@
Following  Kandpal et al. (2023) and  Liu et al. (2024), we use accuracy as our primary evaluation metric, judging whether any of the correct answers appear in the predicted output. To mitigate biases arising from output length, we set a maximum token limit of 200 for the output sequences.
@@
Figure 3: The Inference Pipeline of Block-Attention Model
Prompt Format
@@
The format of input prompt follows  Liu et al. (2024). For retrieved passages, we concatenate them in ascending order of retrieval score. An example is shown in Figure 3.
Base Model & Baselines
@@
We implement the Block-Attention mechanism on two base language models: Llama3-8B-base [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B and Mistral-7B-v3.0555https://huggingface.co/mistralai/Mistral-7B-v0.3. The models, after block fine-tuning, are denoted as Llama3-block and Mistral-block. In addition to these, we construct two strong baselines and two weak baselines.
@@
  • Strong baselines: Llama3-vanilla and Mistral-vanilla. These models are derived from fine-tuning the base language models on the training dataset. They can be considered as the upper bounds for the Block-Attention model. We aim for the Block-Attention model’s performance to approach these as closely as possible.
@@
  • Weak baselines: Llama3-block-wo-ft and Mistral-block-wo-ft. These models transition the attention mechanism of Llama3-vanilla and Mistral-vanilla to a Block-Attention mechanism without any block fine-tuning. the outcomes of these models will represent the lower bounds for the Block-Attention model’s effectiveness, given that they have not undergone any adaptation to Block-Attention during the training phase.
@@
Please note that we have not utilized instruct-based models, such as Llama3-instruct, as our base language model. The reason is our inability to entirely re-implement the fine-tuning processes of these models, which precludes a direct comparison between self-attention and Block-Attention methodologies under identical experimental conditions. Instead, we perform vanilla fine-tuning and block fine-tuning on the same models using the same dataset, ensuring a fair comparison.
@@
Additionally, our idea has some similarities with Ratner et al. (2023). However, their focus is on extending the context window rather than efficient inference. When we directly applied their attention mechanism in the RAG scenario, their accuracy was even lower than Llama3-block-wo-ft, so we did not choose it as a baseline.
Setup
@@
All experiments are done with 8 NVIDIA A800 GPUs, we set the learning $\alpha=2\times 10^{-5}$, batch size $b=64$ and Num. of epochs $n=1$ for all models. We use a linear learning rate warmup for the first 20 steps. Its training is in bfloat16 format, using Deepspeed ZeRO Stage 2, Flash-Attention V2 (for supervised fine-tuning) and SDPA (for block training).
TQA 2wiki NQ HQA Average
Llama3-vanilla 72.8 73.4 55.7 69.5 67.9
Llama3-block-wo-ft 62.7 42.1 43.2 39.6 48.0
Llama3-block 73.0 73.3 56.2 68.5 68.4
Mistral-vanilla 66.6 69.5 46.0 56.4 59.6
Mistral-block-wo-ft 51.9 44.6 32.1 27.8 38.7
Mistral-block 70.7 69.0 50.5 58.6 62.8
@@
Table 1: Accuracy of different models on different datasets. Among them, TQA and 2wiki serve as in-domain test sets, while NQ and HQA are out-of-domain test sets. We have not utilized these latter two datasets for fine-tuning. Furthermore, please note that the ”average” column does not depict a simple arithmetic mean of the preceding four columns, as the number of samples contained in each test set varies.
@@ 3.1 Main Results
From the results in Table 1, we can draw two important conclusions: 1) It is not advisable to directly switch from self-attention to Block-Attention, as it will lead to a sharp drop in the accuracy of the model. 2) However, if we use the Block-Attention mechanism in the fine-tuning stage, then the resultant model has almost the same performance as the self-attention model or is even slightly better on some datasets.
@@
The first conclusion is easy to understand: The model has never seen an input sequence in the Block-Attention manner during the training process. Therefore, the sharp drop makes sense. Next, we will focus on analyzing the second conclusion. From Table 1, we can easily find that the models trained with two attention methods, namely Llama3-vanilla and Llama3-block, not only have comparable effects on in-domain test sets but also have little difference on out-domain test sets. This indicates that in the RAG scenario, it is completely feasible to replace self-attention with Block-Attention, and there will be almost no performance loss.
@@
Moreover, the experimental results on Mistral show some interesting results: the accuracy of Block-Attention models actually slightly outperformed the results of self-attention models on two datasets. It raised a conjecture: is Block-Attention a better attention mechanism than the self-attention mechanism? Intuitively, the semantics of each retrieved passage are mutually independent, and the self-attention method may cause the representation of the passage to be interfered by the context and not accurate enough. However, we still need sufficient experiments to verify the correctness of this assertion, so this is only a conjecture and not our conclusion.
@@
Finally, one may still be interested in knowing exactly how many training steps are needed for the model to adapt to the Block-Attention mechanism. Therefore, we counted the accuracy of the Llama3 model on the four test sets at different fine-tuning steps and plotted it in Figure 4. We find that at the beginning stage of fine-tuning, there is a huge performance difference between the two models. It makes sense because the model needs more training steps to adapt to the Block-Attention manner. After about 400 training steps, the model completely adapts to the new attention mechanism, and the two models show comparable accuracy.
3.2 Efficiency

### v3

@@ Abstract
We introduce Block-Attention, an attention mechanism designed to address the increased inference latency and cost in Retrieval-Augmented Generation (RAG) scenarios. Traditional approaches often encode the entire context. Instead, Block-Attention divides retrieved documents into discrete blocks, with each block independently calculating key-value (KV) states except for the final block. In RAG scenarios, by defining each passage as a block, Block-Attention enables us to reuse the KV states of passages that have been seen before, thereby significantly reducing the latency and the computation overhead during inference. The implementation of Block-Attention involves block segmentation, position re-encoding, and fine-tuning the LLM to adapt to the Block-Attention mechanism. Experiments on four RAG benchmarks demonstrate that after block fine-tuning, the Block-Attention model achieves performance comparable to self-attention models (68.4% vs 67.9% on Llama3) or even superior performance (62.8% vs 59.6% on Mistral). Notably, Block-Attention significantly reduces the time to first token (TTFT) and floating point operations (FLOPs) to a very low level. It only takes 45 ms to output the first token for an input sequence with a total length of 32K. Compared to the self-attention models, the time consumption and corresponding FLOPs are reduced by 98.7% and 99.8%, respectively.  [^1] Codes are available at: https://github.com/TemporaryLoRA/Block-Attention
@@ 1 Introduction
However, everything has two sides, and RAG is no exception. Generally, for each user query, in order to ensure that a passage with the “correct knowledge” is successfully recalled, it is a common practice to retrieve multiple passages—typically between 5 to 30 in most scenarios (Kwiatkowski et al., 2019; Joshi et al., 2017). These passages are then incorporated into the input prompt for the LLM. As a result, the inference efficiency decreases significantly due to the increased sequence length of this extended input prompt. Specifically, the inference latency, measured as the time to first token (TTFT), is considerably higher for a RAG-LLM compared to a non-RAG LLM (Li et al., 2023; Zhu et al., 2024).
Given that the passages in the external databases might have been computed, it is natural to restore their KV cache for fast inference and avoid re-computing these passages. Nonetheless, for autoregressive LLMs, the KV states are inherently context-dependent, which means the KV states for the same passage will vary in different contexts. As a result, when encountering an unseen query, the model must undertake a complete re-encoding of the KV states to ensure an accurate prediction.
In this paper, we propose the Block-Attention method, which reduces the TTFT and computation FLOPs of RAG-LLMs to a level nearly equivalent to that of non-RAG LLMs, while fully maintaining the same accuracy level. As shown in Figure 1, the main idea of Block-Attention is to divide the entire input sequence into several blocks. Each block independently calculates its KV states through self-attention, without any attention to other blocks. Only the final block is able to attend other blocks, i.e., the user query is able to attend all the retrieved documents in the previous blocks. When utilizing Block-Attention in RAG scenarios, we may achieve substantial benefits by defining each passage as a single block and caching its KV states in the memory for further reuse.
@@
The implementation of Block-Attention can be easily achieved through the following steps: 1) Encode all blocks except the last one separately; 2) calculating the positional encoding for each based on its position in the input prompt; 3) Integrating these blocks to compute the KV states for the final block. However, the primary challenge in utilizing Block-Attention is that the LLMs have not been exposed to such an attention mechanism during their training, leading to difficulties in accurately interpreting the input prompt. In our preliminary experiments, we attempted to directly implement Block-Attention in the LLMs without updating any parameters. Unfortunately, this approach resulted in a substantial decrease in performance, with the average accuracy of Llama3-8B on four RAG benchmarks falling from 67.9% to 48.0%.
@@
We evaluate Block-Attention on four RAG benchmarks. Experimental results demonstrate that after fine-tuning, the average accuracy of the Block-Attention model on the benchmarks can remain comparable to (68.4% vs 67.9% on Llama3) or even slightly exceed (62.3% vs 59.6% on Mistral) the performance of self-attention models. In terms of efficiency, we counted the TTFT and FLOPs to the first token (FLOPs-TFT) of the block-attention model when the length of user input is 50 and the total length of the input sequence gradually increases. We found that the longer the total length, the more obvious the improvement of block-attention on efficiency. When the length of the input sequence reaches 32K, the TTFT and FLOPs-TFT of the block-attention model are 1.3% and 0.2% of that of the self-attention model, respectively. [^2] Given that the KV cache is already a mature and low-cost technology (Qin et al., 2024; Lan et al., 2023; Lee et al., 2021), in this paper we do not take the cost of KV cache into account.
@@ 2.1 Main Idea
In the scenario of RAG, a given passage frequently recurs in the retrieval results for numerous queries (Qin et al., 2024), making it a natural consideration to explore the possibility of reusing its KV states. Unfortunately, the auto-regressive nature of the encoding approach, as previously outlined, imposes constraints that render the KV states nearly non-reusable across distinct queries. Our research is dedicated to examining a novel attention mechanism, Block-Attention, that designed to necessitate the re-computation of only the altered text blocks between two input sequences, thereby achieving an outcome equivalent to that of full sequence re-encoding.
@@ 2.3 Position Re-Encoding
2) We rotating $x_{i}$ counterclockwise by $i\theta$ degrees, to re-set its positional encoding to zero:
@@
3) Then, by performing a clockwise rotation of $(i_{\Delta})\theta$ degrees, we obtain the final positional encoding vector:
@@ 2.5 Inference
In inference, the Block-Attention model retrieves the KV states of blocks from the KV cache and concatenates them into the complete input KV states. The detailed process of the inference stage is depicted in Figure 2. Initially, we query and extract the KV states of the first $k-1$ blocks from the cache. Then, based on the position of each block within the input sequence, we calculate their positional encoding using Equation 3. Finally, using the KV states of the first $k-1$ blocks, we compute the KV states of the last block as well as the model output. In the RAG scenarios, the last block is the user query.
@@ 3 Experiments
After the above analysis, there exists two concerns about the Block-Attention method: 1) Can the Block-Attention model achieve the same accuracy as self-attention in RAG scenarios? 2) How much can the Block-Attention mechanism improve the efficiency? The following experimental results will reveal the answers to these two questions. In Sections 3.5, we analyze the accuracy of Block-Attention models. Meanwhile, in Section 3.6, we demonstrate the efficiency of Block-Attention.
3.1 Datasets
Train Dataset
We construct the training datasets by randomly sampling 20,000 instances from these TriviaQA (TQA) (Joshi et al., 2017) and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020) for fine-tuning models. Each training sample consists of (1) a question, (2) 10 passages retrieved from these two datasets using the Contriver toolkit [^3] https://github.com/facebookresearch/contriever, which identifies the 10 most relevant passages, and (3) an answer generated by GPT-4 based on the retrieved passages. The reason to use the GPT-4 answer instead of the ground-truth answers is because the answer might not be present in our retrieved passages. This discrepancy could lead the model to overlook the content of the retrieved passages and generate outputs directly.
Evaluation Dataset
@@ 3.1 Datasets
We evaluate the performance of our proposed Block-Attention mechanism and baseline models on four widely-used RAG benchmarks: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), HotpotQA (HQA) (Yang et al., 2018), and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020). Since the training samples include instances from TQA and 2Wiki, their test sets are considered in-domain, whereas HQA and NQ are treated as out-of-domain test sets. Following  Kandpal et al. (2023) and  Liu et al. (2024), we use accuracy as our primary evaluation metric, judging whether any correct answers appear in the predicted output. To mitigate biases arising from output length, we set a maximum token limit of 200 for the output sequences.
@@
Figure 3: The Inference Pipeline of Block-Attention Model. The retrieved documents at the top have the highest relevance to the user query.
3.2 Input Format
@@
The format of input prompt for all datasets follows  Liu et al. (2024). For retrieved passages, we concatenate them in ascending order of retrieval score. An example is shown in Figure 3.
3.3 Base Model & Baselines
@@
We implement the Block-Attention mechanism on two base language models: Llama3-8B-base [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B and Mistral-7B-v3.0555https://huggingface.co/mistralai/Mistral-7B-v0.3. After applying block fine-tuning, these models are denoted as Llama3-block-ft and Mistral-block-ft. For comparison, we implemented three baselines for each of these two base language models.
@@
  • Strong baselines: Llama3-vanilla-sft and Mistral-vanilla-sft are established by fine-tuning the base language models on our constructed training dataset. These models serve as the upper bound performance for the Block-Attention models. We aim for the Block-Attention model’s performance to approach these as closely as possible.
@@
  • Weak baselines: Llama3-block-w/o-ft and Mistral-block-w/o-ft. These models transition the attention mechanism of Llama3-vanilla-sft and Mistral-vanilla-sft to a Block-Attention mechanism without any block fine-tuning. The outcomes of these models represent the lower bounds for the Block-Attention model’s effectiveness, given that they have not undergone any adaptation to Block-Attention during the training phase.
  • w/o-pos: Llama3-block-ft-w/o-pos and Mistral-block-ft-w/o-pos are baselines that also fine-tuned to adapt the Block-Attention mechanism, while no additional position re-encoding operations described in Section 2.3 are conducted. This group of models will be used to evaluate the effectiveness of the proposed position re-encoding process.
@@
Please note that we have not utilized instruction-tuning models as our base language model, such as Llama3-instruct. The reason is our inability to entirely re-implement the fine-tuning processes of these models, which precludes a direct comparison between self-attention and Block-Attention methodologies under identical experimental conditions. Instead, we perform vanilla fine-tuning and block fine-tuning on the same base models using the same dataset, ensuring a fair comparison.
@@
Additionally, our idea has some similarities with PCW  (Ratner et al., 2023b) and Prompt Cache (Gim et al., 2024). However, the focus of PCW is on extending the context window rather than efficient inference. When we directly applied their attention mechanism in the RAG scenario, their accuracy was even lower than the weak baselines, so we did not choose it as a baseline. We also re-implemented Prompt Cache. Then, we had an interesting finding: their results are exactly the same as the two weak baselines: Llama3-block-w/o-ft and Mistral-block-w/o-ft. Therefore, for the results of Prompt Cache, please directly refer to the result of these two models. [^6] Please note that in addition to accuracy, Prompt Cache also has serious efficiency issues when applied in the RAG field: once the rank of a passage in the input sequence changes, its KV cache cannot be reused.
3.4 Setup
@@
All experiments are done with 8 NVIDIA A800 GPUs, we set the learning rate $\alpha=2\times 10^{-5}$, batch size $b=64$ and the number of epochs $n=1$ for all models. We use a linear learning rate warmup for the first 20 steps. Its training is in bfloat16 format, using DeepSpeed ZeRO Stage-2777https://github.com/microsoft/DeepSpeed and Flash-Attention V2 (Dao et al., 2022). The Flash-Attention toolkit is used for efficient inference for our models and baselines.
Models TQA 2wiki NQ HQA
In-Dom.
Avg.
Out-Dom.
Avg.
Mix
Avg.
Llama3-vanilla-sft 72.8 73.4 55.7 69.5 73.9 62.0 67.9
Llama3-block-ft 73.0 73.3 56.2 68.5 73.6 63.2 68.4
Llama3-block-ft-w/o-pos 70.8 66.8 52.9 61.4 69.3 58.5 63.9
Llama3-block-w/o-ft 62.7 48.1 43.3 39.6 54.7 41.4 48.0
Mistral-vanilla-sft 66.6 69.5 46.0 56.4 68.3 50.9 59.6
Mistral-block-ft 69.9 70.5 50.8 57.9 70.2 54.4 62.3
Mistral-block-ft-w/o-pos 67.8 67.4 47.5 52.4 67.3 50.4 58.8
Mistral-block-w/o-ft 51.9 44.6 32.1 27.8 47.7 29.7 38.7
@@
Table 1: Accuracy of different models on four benchmarks. Please note that the “average” column does not depict a simple arithmetic mean of the preceding columns, as the number of samples contained in each test set varies.
@@ 3.5 Main Results
From the results in Table 1, we can draw three important conclusions:
1) It is not advisable to directly switch from self-attention to Block-Attention, as it will lead to a sharp accuracy drop. For example, removing the Block fine-tune process leads to the 21.99% average absolute performance decrease on all four benchmarks for Llama3-8B and Mistral-7B models.
2) However, if we use the Block-Attention mechanism in the fine-tuning stage, then the resultant model has almost the same performance as the self-attention model or is even slightly better on some datasets. For instance, Mistral-block-ft outperforms the strong baseline Mistral-vanilla-sft, with 2.7% absolute improvements on four benchmarks.
3) The position re-encoding operations are essential for the Block-Attention Model. Removing it leads to significant performance drop—average 4% decreases in accuracy on four datasets.
@@
The first conclusion is easy to understand: The model has never seen an input sequence in the Block-Attention manner during the training process. Therefore, the sharp drop makes sense. Next, we will focus on the second conclusion. From Table 1, we can easily find that the models trained with two attention methods, namely Llama3-vanilla-sft and Llama3-block-ft, not only have comparable effects on in-domain test sets but also have little difference on out-domain test sets. This indicates that in the RAG scenario, it is completely feasible to replace self-attention with Block-Attention, and there will be almost no performance loss.
@@
Moreover, the experimental results on Mistral show some interesting results: the accuracy of Block-Attention models actually slightly outperformed the results of self-attention models on most datasets. It raised a conjecture: is Block-Attention a better attention mechanism than the self-attention mechanism? Intuitively, the semantics of each retrieved passage are mutually independent, and the self-attention method may cause the representation of the passage to be interfered by the context and not accurate enough. However, we still need sufficient experiments to verify the correctness of this assertion, so this is only a conjecture and not our conclusion.
@@
Finally, one may still be interested in knowing exactly how many training steps are needed for the model to adapt to the Block-Attention mechanism. Therefore, we counted the accuracy of the Llama3-8B model on the four test sets at different fine-tuning steps and plotted it in Figure 4. It can be observed that at the beginning stage of fine-tuning, there is a huge performance difference between the two models. It makes sense because the model needs more training steps to adapt to the Block-Attention manner. After about 400 training steps, the model completely adapts to the new attention mechanism, and the two models show comparable accuracy.
3.6 Inference Efficiency of Block-Attention